Using the table below. Answer the following questions: 9.08 8.804:17 267 3.66 6.843.731.63 9.54 3.04
22- If the observations were collected in pairs, the degrecs of freedom when comparing true means of the2 populations would be: A- 4B-5C-6 D-TE-8 23- Calculate the pooled standard deviation: A-3DB-1.76 C-2.96 D-2.93 E- 3.72 Continuing, using α 0.6 and for the following null and alternative hyp following: othesis , Hoi Ha- lg and H : μλ μΒ. Solve the ssumc unknown unequal variances, we can conclude the following about the null hypothesis Ho, A- Accept H>B- Accept Hu
Answer #1

Ans:

A B
1 9.08 6.84
2 8.8 3.73
3 4.17 1.63
4 2.67 9.54
5 3.66 3.04
mean= 5.676 4.956
std. dev.= 3.030 3.194

22)df=5+5-2=8

23)pooled standard deviation=sqrt((4*3.03^2+4*3.194^2)/8)=3.11

24)standard error for difference=3.11*sqrt((1/5)+(1/5))=1.969

Test statistic:

t=(5.676-4.956)/1.969

t=0.366

Fail to reject H0 (i.e. Accept H0)